# Cost-Efficient Long Code Translation using LLMs while Leveraging Identifier Replacements

### v1

@@ 1 Introduction
Our Contributions:
To summarize, the following are our contributions in this work.
@@
    We present a novel zero-shot code translation method that replaces lengthy, complex user-defined identifiers with shorter, generalized placeholders. This approach significantly reduces token count and memory usage, resulting in a more efficient and cost-effective solution for translating long code using LLMs.
@@
    We provide empirical evidence showing that our proposed solution effectively manages long code sequences while preserving both syntactical and hierarchical information during the translation task.
@@ 85 $\triangleright$ Return updated mapping and modified code
We utilize tree-sitter Brunsfeld and contributors , a robust parsing tool, to systematically extract relevant long identifiers from source code. By generating accurate syntax trees across various programming languages, tree-sitter enables us to identify key programming constructs, which are then considered as primary candidates [^2] A detailed statistics of the probable identifier list has been reported in Table 4 for identifier replacements. Following are the key programming constructs serve as primary candidates for identifier extraction and replacement.
@@
Employing these key programming constructs, we developed grammars for each programming language which are then used to extract the potential identifiers suitable for replacements. Let $\mathcal{I}_{s}=\{i_{1},i_{2},\ldots,i_{k}\}$ represent the set of extracted identifiers. For each $i_{j}\in\mathcal{I}_{s}$, we assign a unique placeholder $p_{j}\in\mathcal{P}$ of a small token length = $2$ and produce a simplified source code $C^{\prime}_{s}$ from original source code $C_{s}$. This way, we achieve the length reduction, $\Delta l$, in the token length as given by: $\Delta l=\sum_{j=1}^{k}\left(|i_{j}|-|p_{j}|\right)$, where $|i_{j}|$ and $|p_{j}|$ denote the token length of the original identifier and its corresponding placeholder, respectively. After translation is complete and the LLM generates the output sequence $C^{\prime}_{t}$ in the target language, a reverse process is applied to restore the original identifiers. Let $g$ represent the reverse mapping function, which takes translated sequence $C^{\prime}_{t}$ and original set of identifiers $\mathcal{I}_{t}$ in the target language, and produces the final translated sequence $C_{t}$:
@@
This post-processing step ensures that placeholders in the translated code are replaced with the corresponding identifiers from the target language, preserving the semantics of the original source code. This complete method is described in the Algorithm 1.
@@ 4.1 Dataset
To analyze our approach of code translation via identifier replacement, we partitioned the dataset into three subsets based on token length: those with lengths greater than or equal to 2000, 4000, and 8000 tokens. These lengths were determined using the TikToken tokenizer Pan et al. (2024c). We also filtered out instances that lacked corresponding test cases to ensure accurate evaluation. Table 1 represents the total number of source code instances across various languages for both the original and reduced subsets. The “Original Dataset" provide raw counts prior to filtering, while the “Reduced Dataset" columns illustrate the instances that were retained.
@@ 4.2 LLM Models
In this work, we use models such as GPT3.5-Turbo Ye et al. (2023), GPT4o-Mini Hurst et al. (2024), Mixtral (8*7B) Jiang et al. (2023) [^3] It is noteworthy that, while CodeLlama and Mixtral models are theoretically capable of handling inputs more than the mentioned tokens, they often produce noise beyond this limit., and CodeLlama Grattafiori et al. (2023). The context window of these models during inference are 16K, 128K, 8K and 4K respectively. The model choice is driven by their widespread use and popularity, as well as the constraints of computing capabilities.
@@ 6 Ablation Study
Table 5: Performance comparison of Java and C++ translations across different identifier replacement categories. The table evaluates the translation performance when selectively replacing different identifier types (Declaration, Function, Parameter, and Expression) for Java and C++ while translating them into C, C++, Go, and Python. The accuracy (Acc.) represents the mean accuracy across the target languages for each identifier category, while the IdRep Accuracy (IdRep Acc.) represents the average accuracy when all identifier types are replaced and the NoRep Accuracy (NoRep Acc.) represents the average accuracy without any replacement as reported in Table 2. The best performance is in bold faced.
@@ 7 Conclusion and Future Directions
We present a novel zero-shot approach for long code translation that leverages identifier replacement to overcome the context window and memory limitations of LLMs. By substituting long identifiers with shorter placeholders, the method reduces token count, leading to cost-effective and efficient translation without sacrificing code integrity. Our experiments demonstrate significant cost savings and improved translation accuracy.
Future research can focus on hybrid techniques, such as combining identifier replacement with symbolic execution, to further boost translation performance.

### v2

# LLM Based Long Code Translation using Identifier Replacement

@@ 1 Introduction
The following are our contributions in this work.
@@
    We introduce a novel zero-shot code translation technique that substitutes lengthy, complex user-defined identifiers with concise, generalized placeholders. This method substantially decreases token count and memory consumption, leading to a more efficient and cost-effective process for translating long code with LLMs.
@@
    The empirical evidence demonstrates that our proposed solution effectively handles long code sequences while maintaining both syntactical and hierarchical information during translation.
@@ 85 $\triangleright$ Return updated mapping and modified code
In this method as described in the Algorithm 1, we employ tree-sitterBrunsfeld and contributors , a powerful parsing tool, to systematically extract relevant long identifiers from source code. By generating precise syntax trees across multiple programming languages, tree-sitter allows us to identify key programming constructs, which serve as primary candidates [^2] A detailed statistical overview of the probable identifier list is provided in Table4 for identifier replacement. The following key programming constructs are the main candidates used for identifier extraction and replacement.
@@
Employing these key programming constructs, we developed grammars for each programming language which are then used to extract the potential identifiers suitable for replacements. Let $\mathcal{I}_{s}=\{i_{1},i_{2},\ldots,i_{k}\}$ represent the set of extracted identifiers. For each $i_{j}\in\mathcal{I}_{s}$, we assign a unique placeholder $p_{j}\in\mathcal{P}$ of a small token length = $2$ and produce a simplified source code $C^{\prime}_{s}$ from original source code $C_{s}$. This way, we achieve the length reduction, $\Delta l$, in the token length as given by: $\Delta l=\sum_{j=1}^{k}\left(|i_{j}|-|p_{j}|\right)$, where $|i_{j}|$ and $|p_{j}|$ denote the token length of the original identifier and its corresponding placeholder, respectively.
After translation is complete and the LLM generates the output sequence $C^{\prime}_{t}$ in the target language, a reverse process is applied to restore the original identifiers. Let $g$ represent the reverse mapping function, which takes translated sequence $C^{\prime}_{t}$ and original set of identifiers $\mathcal{I}_{t}$ in the target language, and produces the final translated sequence $C_{t}$:
@@
This post-processing step ensures that placeholders in the translated code are replaced with the corresponding identifiers from the target language, preserving the semantics of the original source code.
@@ 4.1 Dataset
To analyze our approach of code translation via identifier replacement, we partitioned the dataset into three subsets based on token length: those with lengths greater than or equal to 2000, 4000, and 8000 tokens. These lengths were determined using the TikToken tokenizer Pan et al. (2024c). We also excluded instances without corresponding test cases to ensure accurate assessment. Table1 presents the total number of source code instances across various languages for both the original and reduced subsets. The "Original Dataset" columns show raw counts before filtering, while the "Reduced Dataset" columns indicate the instances retained after filtering.
@@ 4.2 LLM Models
In this work, we employ models such as GPT-3.5-Turbo Ye et al. (2023), GPT4o-Mini Hurst et al. (2024), Mixtral (8×7B) Jiang et al. (2023), and CodeLlama Grattafiori et al. (2023). The choice of these models is primarily driven by their widespread adoption, popularity, and the constraints of available computational resources. Although some of these models can theoretically handle inputs exceeding the token limits of the dataset mentioned earlier, we can still assess the effectiveness of our identifier replacement technique for long code translation tasks.
@@ 6 Ablation Study
Table 5: Performance comparison of Java and C++ translations across different identifier replacement categories. The table evaluates the translation performance when selectively replacing different identifier types (Declaration, Function, Parameter, and Expression) for Java and C++ while translating them into C, C++, Go, and Python. The accuracy (Acc.) represents the mean accuracy across the target languages for each identifier category, while the IdRep Accuracy (IdRep Acc.) represents the average accuracy when all identifier types are replaced and the NoRep Accuracy (NoRep Acc.) represents the average accuracy without any replacement as reported in Table 2. The best performance is in bold.
@@ 7 Conclusion and Future Directions
We present a novel zero-shot approach for long code translation that leverages identifier replacement to overcome the context window and memory limitations of LLMs. By substituting long identifiers with shorter placeholders, the method reduces token count, leading to cost-effective and efficient translation without sacrificing code integrity. Our experiments demonstrate significant cost savings and improved translation accuracy. Future research can focus on hybrid techniques, such as combining identifier replacement with symbolic execution, to further boost translation performance.